# An improved formula for Wigner function and spin polarization in a decoupling relativistic fluid at local thermodynamic equilibrium

### v2

@@ Acknowledgments.
$\widehat{T}^{\mu\nu}(x)=\frac{i}{4}\bar{\psi}(x)\left[\gamma^{\mu}\left(\overrightarrow{\partial}^{\nu}-\overleftarrow{\partial}^{\nu}\right)+\gamma^{\nu}\left(\overrightarrow{\partial}^{\mu}-\overleftarrow{\partial}^{\mu}\right)\right]\psi(x)\,$, (46)
Using the quantized field operator in eq. (24), we can express $\widehat{T}^{\mu\nu}(0)$ as (with $p=(p_{+}+p_{-})/2$):
$\widehat{T}^{\mu\nu}(0)=\frac{1}{2(2\pi)^{3}}\sum_{r,s}\int{\rm d}^{4}p_{+}{\rm d}^{4}p_{-}\;\delta(p_{+}^{2}-m^{2})\delta(p_{-}^{2}-m^{2})\bar{u}_{s}(p_{-})(\gamma^{\mu}p^{\nu}+\gamma^{\nu}p^{\mu})u_{r}(p_{+})\widehat{a}_{s}^{\dagger}(p_{-})\widehat{a}_{r}(p_{+})\theta(p^{0}_{+})\theta(p^{0}_{-})$ (47)
@@
$\left\langle\widehat{a}_{r}^{\dagger}(p_{+})\widehat{a}_{s}(p_{-}),\widehat{T}^{\mu\nu}(0)\right\rangle_{c,\text{GE}}$ $=$ $\frac{1}{2(2\pi)^{3}}\int{\rm d}^{4}p_{+}^{\prime}{\rm d}^{4}p_{-}^{\prime}\,\delta(p_{+}^{\prime 2}-m^{2})\delta(p_{-}^{\prime 2}-m^{2})\theta(p^{\prime 0}_{+})\theta(p^{\prime 0}_{-})$ (48)
@@
with $\varepsilon_{\pm}=\sqrt{m^{2}+{\bf p}_{\pm}^{2}}$ being the on-shell energies. Substituting the equation (48) into (Wigner function.) and using the relation:
@@
$\text{tr}\left[\gamma^{\mu}\gamma^{5}G^{\nu\alpha}(p,q)\right]=-i\frac{n_{F}(x,p_{+})-n_{F}(x,p_{-})}{(2\pi)^{3}(\beta\cdot q)}\delta\left(p^{2}+\frac{q^{2}}{4}-m^{2}\right)\theta(p_{+}^{0})\theta(p_{-}^{0})(g_{\rho}^{\nu}p^{\alpha}+g_{\rho}^{\alpha}p^{\nu})\epsilon^{\mu\rho\lambda\tau}p_{\lambda}q_{\tau}\,$, (49)
@@
${\cal A}_{\text{LE}}^{\mu}(x,p)$ $=$ $-i\sum_{N=0}^{\infty}\sum_{\bar{y}(x,p)}\frac{1}{N!}D^{N}_{y}\left[\frac{n_{F}(x,p_{+})-n_{F}(x,p_{-})}{(2\pi)^{3}(\beta\cdot q)}\delta\left(p^{2}+\frac{q^{2}}{4}-m^{2}\right)\theta(p_{+}^{0})\theta(p_{-}^{0})\right$. (50)
$\left.\times(g_{\rho}^{\nu}p^{\alpha}+g_{\rho}^{\alpha}p^{\nu})\epsilon^{\mu\rho\lambda\tau}p_{\lambda}q_{\tau}\,\frac{n_{\mu}(y)}{|p\cdot n(y)|}\Delta\beta_{\nu}(y,x)\right]\Bigg|_{q=0,y=\bar{y}(x,p)}\,.$
@@
$\mathcal{A}^{\mu}_{\text{LE}}(x,p)$ $\simeq$ $\frac{2}{(2\pi)^{3}}n_{F}(x,p)[1-n_{F}(x,p)]\delta(p^{2}-m^{2})\epsilon^{\mu\rho\lambda\tau}p_{\lambda}$ (51)
@@
$p\cdot n(x)]\frac{1}{2}\epsilon^{\mu\nu\rho\lambda}p_{\nu}\left[\left.\partial_{\rho}^{y}\beta_{\lambda}(y)\right|_{y=\bar{y}(x,p)}\right]-\frac{1}{2}\epsilon^{\mu\nu\rho\lambda}p_{\nu}n_{\rho}(x)p^{\alpha}\left[\left.\partial_{\lambda}^{y}\beta_{\alpha}(y)+\partial_{\alpha}^{y}\beta_{\lambda}(y)\right|_{y=\bar{y}(x,p)}\right]$ (52)
@@
in the eq. (52) and reinstating the missing factors, we obtain the first part of $\mathcal{A}_{\text{LE}}^{\mu}$:
$\mathcal{A}_{\text{LE}-1}^{\mu}(x,p)=-\frac{\delta(p^{2}-m^{2})}{(2\pi)^{3}}n_{F}(x,p)(1-n_{F}(x,p))\sum_{\bar{y}(x,p)}\text{sgn}[p\cdot n(\bar{y})]\epsilon^{\mu\nu\rho\lambda}p_{\nu}\left[\varpi_{\rho\lambda}(\bar{y})+\frac{2}{p\cdot n(\bar{y})}n_{\rho}(\bar{y})\xi_{\lambda\alpha}(\bar{y})p^{\alpha}\right]\,.$ (53)
@@
$\epsilon^{\mu\rho\lambda\tau}p_{\lambda}(g_{\rho}^{\nu}p^{\alpha}+g_{\rho}^{\alpha}p^{\nu})\Delta_{\tau\kappa}(\bar{y})\Delta\beta_{\alpha}(\bar{y},x)\left.\left[\partial_{y}^{\kappa}\frac{n_{\nu}(y)}{|p\cdot n(y)|}\right]\right|_{y=\bar{y}}\,$, (54)
@@
$\epsilon^{\mu\rho\lambda\tau}p_{\lambda}(g_{\rho}^{\nu}p^{\alpha}-g_{\rho}^{\alpha}p^{\nu})\Delta_{\tau\kappa}(\bar{y})\Delta\beta_{\alpha}(\bar{y},x)\left.\left[\partial_{y}^{\kappa}\frac{n_{\nu}(y)}{|p\cdot n(y)|}\right]\right|_{y=\bar{y}}\,$, (55)
@@
$\epsilon^{\mu\nu\lambda\tau}p^{\alpha}+\epsilon^{\nu\lambda\tau\alpha}p^{\mu}+\epsilon^{\lambda\tau\alpha\mu}p^{\nu}+\epsilon^{\tau\alpha\mu\nu}p^{\lambda}+\epsilon^{\alpha\mu\nu\lambda}p^{\tau}=0\,$, (56)
@@
Plugging this relation into eq. (55) and taking into account that $p^{\tau}\Delta_{\tau\kappa}(\bar{y})=0$, the expression in (55) can be rewritten as:
$-(\epsilon^{\nu\lambda\tau\alpha}p^{\mu}+\epsilon^{\tau\alpha\mu\nu}p^{\lambda})p_{\lambda}\Delta_{\tau\kappa}(\bar{y})\Delta\beta_{\alpha}(\bar{y},x)\left.\left[\partial_{y}^{\kappa}\frac{n_{\nu}(y)}{|p\cdot n(y)|}\right]\right|_{y=\bar{y}}$ (57)
@@
$\epsilon^{\nu\lambda\tau\alpha}\Delta_{\tau}^{\ \kappa}(\bar{y})$ $=$ $\epsilon^{\nu\lambda\tau\alpha}\left[g_{\tau}^{\kappa}-\frac{p^{\kappa}n_{\tau}(\bar{y})}{p\cdot n(\bar{y})}\right]=\left(p^{\tau}\epsilon^{\nu\lambda\kappa\alpha}-p^{\kappa}\epsilon^{\nu\lambda\tau\alpha}\right)\frac{n_{\tau}(\bar{y})}{p\cdot n(\bar{y})}$ (58)
@@
where we have again used the Schouten identity (56) in the last step. Substituting (58) into eq. (57) and noting that (see also above discussion):
@@
the expression in eq. (57) gets its final form:
$(p^{\mu}p_{\lambda}-g_{\lambda}^{\mu}p^{2})\epsilon^{\lambda\tau\nu\kappa}\frac{n_{\tau}(\bar{y})}{p\cdot n(\bar{y})}\left.\left[\partial_{\kappa}^{y}\frac{n_{\nu}(y)}{|p\cdot n(y)|}\right]\right|_{y=\bar{y}}p^{\alpha}\Delta\beta_{\alpha}(\bar{y},x)\,.$ (59)
@@
$\Omega^{\lambda}(\bar{y})\equiv\epsilon^{\lambda\tau\nu\kappa}\frac{n_{\tau}(\bar{y})}{p\cdot n(\bar{y})}\left.\left[\partial_{\kappa}^{y}\frac{n_{\nu}(y)}{|p\cdot n(y)|}\right]\right|_{y=\bar{y}}=\frac{\text{sgn}[p\cdot n(\bar{y})]}{[p\cdot n(\bar{y})]^{2}}\epsilon^{\lambda\tau\nu\kappa}n_{\tau}(\bar{y})\left.\left[\partial_{\kappa}^{y}n_{\nu}(y)\right]\right|_{y=\bar{y}}\,$, (60)
which depends on the curvature of the decoupling hypersurface at point $\bar{y}$, the equation (59) can be rewritten as:
$(p^{\mu}p_{\lambda}-g_{\lambda}^{\mu}p^{2})p^{\alpha}\Delta\beta_{\alpha}(\bar{y},x)\Omega^{\lambda}(\bar{y})\,.$ (61)
@@
$\partial_{\kappa}n_{\nu}=A\,\partial_{\kappa}\sigma_{\nu}+(\partial_{\kappa}A)n_{\nu}$ (62)
@@
it turns out that the gradient of $\sigma$ in eq. (62) is symmetric under the exchange of its indices. Therefore, the contraction with $\epsilon^{\lambda\tau\nu\kappa}$ in equation (60) of both the terms in equation (62) vanishes and so $\Omega^{\lambda}=0$. In conclusion, the expression in (59) vanishes, demonstrating that the curvature of $\Sigma_{\rm D}$ does not contribute to the spin polarization at the leading order.
@@
Finally, plugging eqs. (45) and (53) into eq. (22), the formula (23) is obtained.
S4 Isothermal decoupling hypersurface
@@
$\xi^{\mu\nu}$ $=$ $\frac{1}{2T}(\partial^{\mu}u^{\nu}+\partial^{\nu}u^{\mu})-\frac{1}{2T^{2}}(u^{\nu}\partial^{\mu}T+u^{\mu}\partial^{\nu}T)\,.$ (63)
Substituting them into eq. (23), the vorticity-induced polarization and shear-induced polarization read:
@@
$S_{\text{shear}}^{\mu}(p)$ $=$ $-\frac{1}{8mN_{p}}\int_{\Sigma_{\rm D}}\!\!{\rm d}\Sigma(x)\cdot p\,n_{F}(x,p)(1-n_{F}(x,p))\epsilon^{\mu\nu\rho\lambda}p_{\nu}$ (64)
@@
$S_{\text{T-grad}}^{\mu}(p)$ $=$ $-\frac{1}{8mN_{p}}\int_{\Sigma_{\rm D}}\!\!{\rm d}\Sigma(x)\cdot p\,n_{F}(x,p)[1-n_{F}(x,p)]\epsilon^{\mu\nu\rho\lambda}p_{\nu}$ (65)
$\times\sum_{\bar{y}(x,p)}\text{sgn}[p\cdot n]\frac{1}{T^{2}}\left\{u_{\lambda}\left[\partial_{\rho}T-n_{\rho}\frac{p\cdot\partial T}{p\cdot n}\right]-\frac{p\cdot u}{p\cdot n}n_{\rho}\partial_{\lambda}T]\right\}\Bigg|_{\bar{y}(x,p)}\,.$

### v3

@@ Acknowledgments.
$\widehat{T}^{\mu\nu}(x)=\frac{i}{4}\bar{\psi}(x)\left[\gamma^{\mu}\left(\overrightarrow{\partial}^{\nu}-\overleftarrow{\partial}^{\nu}\right)+\gamma^{\nu}\left(\overrightarrow{\partial}^{\mu}-\overleftarrow{\partial}^{\mu}\right)\right]\psi(x)\,$, (S23)
Using the quantized field operator in eq. (S1), we can express $\widehat{T}^{\mu\nu}(0)$ as (with $p=(p_{+}+p_{-})/2$):
$\widehat{T}^{\mu\nu}(0)=\frac{1}{2(2\pi)^{3}}\sum_{r,s}\int{\rm d}^{4}p_{+}{\rm d}^{4}p_{-}\;\delta(p_{+}^{2}-m^{2})\delta(p_{-}^{2}-m^{2})\bar{u}_{s}(p_{-})(\gamma^{\mu}p^{\nu}+\gamma^{\nu}p^{\mu})u_{r}(p_{+})\widehat{a}_{s}^{\dagger}(p_{-})\widehat{a}_{r}(p_{+})\theta(p^{0}_{+})\theta(p^{0}_{-})$ (S24)
@@
$\left\langle\widehat{a}_{r}^{\dagger}(p_{+})\widehat{a}_{s}(p_{-}),\widehat{T}^{\mu\nu}(0)\right\rangle_{c,\text{GE}}$ $=$ $\frac{1}{2(2\pi)^{3}}\int{\rm d}^{4}p_{+}^{\prime}{\rm d}^{4}p_{-}^{\prime}\,\delta(p_{+}^{\prime 2}-m^{2})\delta(p_{-}^{\prime 2}-m^{2})\theta(p^{\prime 0}_{+})\theta(p^{\prime 0}_{-})$ (S25)
@@
with $\varepsilon_{\pm}=\sqrt{m^{2}+{\bf p}_{\pm}^{2}}$ being the on-shell energies. Substituting the equation (S25) into (Wigner function.) and using the relation:
@@
$\text{tr}\left[\gamma^{\mu}\gamma^{5}G^{\nu\alpha}(p,q)\right]=-i\frac{n_{F}(x,p_{+})-n_{F}(x,p_{-})}{(2\pi)^{3}(\beta\cdot q)}\delta\left(p^{2}+\frac{q^{2}}{4}-m^{2}\right)\theta(p_{+}^{0})\theta(p_{-}^{0})(g_{\rho}^{\nu}p^{\alpha}+g_{\rho}^{\alpha}p^{\nu})\epsilon^{\mu\rho\lambda\tau}p_{\lambda}q_{\tau}\,$, (S26)
@@
${\cal A}_{\text{LE}}^{\mu}(x,p)$ $=$ $-i\sum_{N=0}^{\infty}\sum_{\bar{y}(x,p)}\frac{1}{N!}D^{N}_{y}\left[\frac{n_{F}(x,p_{+})-n_{F}(x,p_{-})}{(2\pi)^{3}(\beta\cdot q)}\delta\left(p^{2}+\frac{q^{2}}{4}-m^{2}\right)\theta(p_{+}^{0})\theta(p_{-}^{0})\right$. (S27)
$\left.\times(g_{\rho}^{\nu}p^{\alpha}+g_{\rho}^{\alpha}p^{\nu})\epsilon^{\mu\rho\lambda\tau}p_{\lambda}q_{\tau}\,\frac{n_{\nu}(y)}{|p\cdot n(y)|}\Delta\beta_{\alpha}(y,x)\right]\Bigg|_{q=0,y=\bar{y}(x,p)}\,.$
@@
$\mathcal{A}^{\mu}_{\text{LE}}(x,p)$ $\simeq$ $\frac{2}{(2\pi)^{3}}n_{F}(x,p)[1-n_{F}(x,p)]\delta(p^{2}-m^{2})\epsilon^{\mu\rho\lambda\tau}p_{\lambda}$ (S28)
@@
$p\cdot n(x)]\frac{1}{2}\epsilon^{\mu\nu\rho\lambda}p_{\nu}\left[\left.\partial_{\rho}^{y}\beta_{\lambda}(y)\right|_{y=\bar{y}(x,p)}\right]-\frac{1}{2}\epsilon^{\mu\nu\rho\lambda}p_{\nu}n_{\rho}(x)p^{\alpha}\left[\left.\partial_{\lambda}^{y}\beta_{\alpha}(y)+\partial_{\alpha}^{y}\beta_{\lambda}(y)\right|_{y=\bar{y}(x,p)}\right]$ (S29)
@@
in the eq. (S29) and reinstating the missing factors, we obtain the first part of $\mathcal{A}_{\text{LE}}^{\mu}$:
$\mathcal{A}_{\text{LE}-1}^{\mu}(x,p)=-\frac{\delta(p^{2}-m^{2})}{(2\pi)^{3}}n_{F}(x,p)(1-n_{F}(x,p))\sum_{\bar{y}(x,p)}\text{sgn}[p\cdot n(\bar{y})]\epsilon^{\mu\nu\rho\lambda}p_{\nu}\left[\varpi_{\rho\lambda}(\bar{y})+\frac{2}{p\cdot n(\bar{y})}n_{\rho}(\bar{y})\xi_{\lambda\alpha}(\bar{y})p^{\alpha}\right]\,.$ (S30)
@@
$\epsilon^{\mu\rho\lambda\tau}p_{\lambda}(g_{\rho}^{\nu}p^{\alpha}+g_{\rho}^{\alpha}p^{\nu})\Delta_{\tau\kappa}(\bar{y})\Delta\beta_{\alpha}(\bar{y},x)\left.\left[\partial_{y}^{\kappa}\frac{n_{\nu}(y)}{|p\cdot n(y)|}\right]\right|_{y=\bar{y}}\,$, (S31)
@@
$\epsilon^{\mu\rho\lambda\tau}p_{\lambda}(g_{\rho}^{\nu}p^{\alpha}-g_{\rho}^{\alpha}p^{\nu})\Delta_{\tau\kappa}(\bar{y})\Delta\beta_{\alpha}(\bar{y},x)\left.\left[\partial_{y}^{\kappa}\frac{n_{\nu}(y)}{|p\cdot n(y)|}\right]\right|_{y=\bar{y}}\,$, (S32)
@@
$\epsilon^{\mu\nu\lambda\tau}p^{\alpha}+\epsilon^{\nu\lambda\tau\alpha}p^{\mu}+\epsilon^{\lambda\tau\alpha\mu}p^{\nu}+\epsilon^{\tau\alpha\mu\nu}p^{\lambda}+\epsilon^{\alpha\mu\nu\lambda}p^{\tau}=0\,$, (S33)
@@
Plugging this relation into eq. (S32) and taking into account that $p^{\tau}\Delta_{\tau\kappa}(\bar{y})=0$, the expression in (S32) can be rewritten as:
$-(\epsilon^{\nu\lambda\tau\alpha}p^{\mu}+\epsilon^{\tau\alpha\mu\nu}p^{\lambda})p_{\lambda}\Delta_{\tau\kappa}(\bar{y})\Delta\beta_{\alpha}(\bar{y},x)\left.\left[\partial_{y}^{\kappa}\frac{n_{\nu}(y)}{|p\cdot n(y)|}\right]\right|_{y=\bar{y}}$ (S34)
@@
$\epsilon^{\nu\lambda\tau\alpha}\Delta_{\tau}^{\ \kappa}(\bar{y})$ $=$ $\epsilon^{\nu\lambda\tau\alpha}\left[g_{\tau}^{\kappa}-\frac{p^{\kappa}n_{\tau}(\bar{y})}{p\cdot n(\bar{y})}\right]=\left(p^{\tau}\epsilon^{\nu\lambda\kappa\alpha}-p^{\kappa}\epsilon^{\nu\lambda\tau\alpha}\right)\frac{n_{\tau}(\bar{y})}{p\cdot n(\bar{y})}$ (S35)
@@
where we have again used the Schouten identity (S33) in the last step. Substituting (S35) into eq. (S34) and noting that (see also above discussion):
@@
the expression in eq. (S34) gets its final form:
$(p^{\mu}p_{\lambda}-g_{\lambda}^{\mu}p^{2})\epsilon^{\lambda\tau\nu\kappa}\frac{n_{\tau}(\bar{y})}{p\cdot n(\bar{y})}\left.\left[\partial_{\kappa}^{y}\frac{n_{\nu}(y)}{|p\cdot n(y)|}\right]\right|_{y=\bar{y}}p^{\alpha}\Delta\beta_{\alpha}(\bar{y},x)\,.$ (S36)
@@
$\Omega^{\lambda}(\bar{y})\equiv\epsilon^{\lambda\tau\nu\kappa}\frac{n_{\tau}(\bar{y})}{p\cdot n(\bar{y})}\left.\left[\partial_{\kappa}^{y}\frac{n_{\nu}(y)}{|p\cdot n(y)|}\right]\right|_{y=\bar{y}}=\frac{\text{sgn}[p\cdot n(\bar{y})]}{[p\cdot n(\bar{y})]^{2}}\epsilon^{\lambda\tau\nu\kappa}n_{\tau}(\bar{y})\left.\left[\partial_{\kappa}^{y}n_{\nu}(y)\right]\right|_{y=\bar{y}}\,$, (S37)
which depends on the curvature of the decoupling hypersurface at point $\bar{y}$, the equation (S36) can be rewritten as:
$(p^{\mu}p_{\lambda}-g_{\lambda}^{\mu}p^{2})p^{\alpha}\Delta\beta_{\alpha}(\bar{y},x)\Omega^{\lambda}(\bar{y})\,.$ (S38)
@@
$\partial_{\kappa}n_{\nu}=A\,\partial_{\kappa}\sigma_{\nu}+(\partial_{\kappa}A)n_{\nu}$ (S39)
@@
it turns out that the gradient of $\sigma$ in eq. (S39) is symmetric under the exchange of its indices. Therefore, the contraction with $\epsilon^{\lambda\tau\nu\kappa}$ in equation (S37) of both the terms in equation (S39) vanishes and so $\Omega^{\lambda}=0$. In conclusion, the expression in (S36) vanishes, demonstrating that the curvature of $\Sigma_{\rm D}$ does not contribute to the spin polarization at the leading order.
@@
Finally, plugging eqs. (S22) and (S30) into eq. (22), the vorticity- and shear-induced polarization in (23) are obtained.
S4 Finite chemical potential
In this section we show the calculations for a system with finite chemical potential. This case applies to intermediate- and low-energy collisions. The density operator takes the following form:
$\widehat{\rho}=\frac{1}{Z}\exp\left[-\int_{\Sigma_{0}}\!\!{\rm d}\Sigma_{\mu}(y)\;\widehat{T}^{\mu\nu}(y)\beta_{\nu}(y)+\int_{\Sigma_{0}}\!\!{\rm d}\Sigma_{\mu}(y)\;\widehat{j}^{\mu}(y)\zeta(y)\right]\,$, (S40)
where $\widehat{j}^{\mu}$ is the conserved current operator, and $\zeta\equiv\mu/T$ is the chemical potential over temperature. Similar to eq. (2) in the main text, the second term in the exponent in eq. (S40) can be transformed into the sum of two terms:
$\int_{\Sigma_{0}}\!\!{\rm d}\Sigma_{\mu}(y)\;\widehat{j}^{\mu}(y)\zeta(y)=\int_{\Sigma_{\rm D}}\!\!{\rm d}\Sigma_{\mu}(y)\;\widehat{j}^{\mu}(y)\zeta(y)+\int_{\Omega}d^{4}y\;\widehat{j}^{\mu}(y)\partial_{\mu}\zeta(y)\,$, (S41)
where the second terms accounts for the dissipative effects induced by the gradients of $\zeta$. For an arbitrary operator $\widehat{O}(x)$, its mean value at LTE is given by:
$O_{\rm LE}(x)=\frac{1}{Z}{\rm Tr}\left(\exp\left[-\int_{\Sigma_{\rm D}}\!\!{\rm d}\Sigma_{\mu}(y)\;\widehat{T}^{\mu\nu}(y)\beta_{\nu}(y)+\int_{\Sigma_{\rm D}}\!\!{\rm d}\Sigma_{\mu}(y)\;\widehat{j}^{\mu}(y)\zeta(y)\right]\right)\,.$ (S42)
We replace $\beta_{\nu}(y)$ in the above equation with $\beta_{\nu}(x)+\Delta\beta_{\nu}(y,x)$ and similarly replace $\zeta(y)$ with:
$\zeta(y)=\zeta(x)+\left(\zeta(y)-\zeta(x)\right)\equiv\zeta(x)+\Delta\zeta(y,x)\,.$ (S43)
Using the linear response theory, the mean value of $\widehat{O}(x)$ at LTE reads:
$O_{\rm LE}(x)=O_{\rm GE}(x)+\Delta O_{\rm LE}(x)+\Delta O_{\rm LE}^{\zeta}(x)\,$, (S44)
where $\Delta O_{\rm LE}$ and $\Delta O_{\rm LE}^{\zeta}$ are linear responses to $\Delta\beta_{\nu}$ and $\Delta\zeta$, respectively. The term $\Delta O_{\rm LE}$ has been discussed in the main text and thus in this section we only focus on $\Delta O_{\rm LE}^{\zeta}$:
$\Delta O_{\rm LE}^{\zeta}(x)=-\int_{\Sigma_{\rm D}}\!\!{\rm d}\Sigma_{\mu}(y)\;\int_{0}^{1}{\rm d}z\;\Delta\zeta(y,x)\left\langle\widehat{O}(x),\,{\rm e}^{-z\beta(x)\cdot{\widehat{P}}}\widehat{j}^{\mu}(y){\rm e}^{z\beta(x)\cdot{\widehat{P}}}\right\rangle_{c,\text{GE}}\,.$ (S45)
Replacing $\widehat{O}$ with the Wigner operator, the linear response to $\Delta\zeta$ can be expressed as:
$\Delta W_{\rm LE}^{\zeta}(x,p)=\frac{1}{(2\pi)^{3}}\int d^{4}q\;\delta(p\cdot q)H^{\mu}(q)\int_{\Sigma_{\rm D}}\!\!{\rm d}\Sigma_{\mu}(y)\;e^{-iq\cdot(y-x)}\Delta\zeta(y,x)\,$, (S46)
where:
$H^{\mu}(q)=\frac{1}{2}\delta\left(p^{2}+\frac{q^{2}}{4}-m^{2}\right)\frac{1-{\rm e}^{\beta(x)\cdot q}}{\beta(x)\cdot q}\theta(p_{+}^{0})\theta(p_{-}^{0})\sum_{r,s}u_{s}(p_{-})\bar{u}_{r}(p_{+})\left\langle\widehat{a}^{\dagger}_{r}(p_{+})\widehat{a}_{s}(p_{-}),\widehat{j}^{\mu}(0)\right\rangle_{c,{\rm GE}}\,.$ (S47)
Similar to the discussion below eq. (9), $\Delta\zeta$ is slowly varying in space and time in the hydrodynamic limit, allowing us to take a small-$q$ expansion of $H^{\mu}(q)$. The results can be obtained from eq. (10) by replacing $G^{\mu\nu}$ with $H^{\mu}$, and $\Delta\beta_{\nu}$ with $\Delta\zeta$:
$\Delta W_{\rm LE}^{\zeta+}(x,p)=\frac{1}{(2\pi)^{3}}\sum_{N=0}^{\infty}\frac{1}{N!}\int_{\Sigma_{\rm D}}\!\!{\rm d}\Sigma_{\mu}(y)\,I_{n}^{\nu_{1}\nu_{2}\cdots\nu_{N}}(y-x)\left.\Delta\zeta(y,x)\left[\partial_{\nu_{1}}^{q}\partial_{\nu_{2}}^{q}\cdots\partial_{\nu}^{q}H^{\mu}(q)\right]\right|_{q=0}\,.$ (S48)
Completing the integral over $\Sigma_{\rm D}$, a compact form can be achieved:
$\Delta W_{\rm LE}^{\zeta+}(x,p)=\sum_{N=0}^{\infty}\sum_{\bar{y}(x,p)}\frac{1}{N!}D^{N}_{y}(\bar{y})\left[H^{\mu}(q)\frac{n_{\mu}(y)}{|p\cdot n(y)|}\Delta\zeta(y,x)\right]\Bigg|_{q=0,y=\bar{y}(x,p)}\,.$ (S49)
Substituting the current operator of free fermions into eq. (S47) we obtain:
[…]
The axial-vector component of $H^{\mu}$ can be evaluated as:
$\text{tr}\left[\gamma^{\mu}\gamma^{5}H^{\nu}(p,q)\right]=i\frac{2\left[n_{F}(x,p_{+})-n_{F}(x,p_{-})\right]}{(2\pi)^{3}(\beta\cdot q)}\delta\left(p^{2}+\frac{q^{2}}{4}-m^{2}\right)\theta(p_{+}^{0})\theta(p_{-}^{0})\epsilon^{\mu\nu\lambda\tau}p_{\lambda}q_{\tau}\,$, (S51)
Then the axial-vector component of the Wigner function has the following linear response to $\Delta\zeta$:
${\cal A}_{\text{LE},\zeta}^{\mu}(x,p)$ $=$ $2i\sum_{N=0}^{\infty}\sum_{\bar{y}(x,p)}\frac{1}{N!}D^{N}_{y}\left[\frac{n_{F}(x,p_{+})-n_{F}(x,p_{-})}{(2\pi)^{3}(\beta\cdot q)}\right$. (S52)
$\left.\times\delta\left(p^{2}+\frac{q^{2}}{4}-m^{2}\right)\theta(p_{+}^{0})\theta(p_{-}^{0})\epsilon^{\mu\nu\lambda\tau}p_{\lambda}q_{\tau}\,\frac{n_{\nu}(y)}{|p\cdot n(y)|}\Delta\zeta(y,x)\right]\Bigg|_{q=0,y=\bar{y}(x,p)}\,.$
The leading order contribution arises from the term of $N=1$, which is given by:
${\cal A}_{\text{LE},\zeta}^{\mu}(x,p)$ $=$ $-\frac{2}{(2\pi)^{3}}\sum_{\bar{y}(x,p)}n_{F}(x,p)[1-n_{F}(x,p)]\delta(p^{2}-m^{2})$ (S53)
$\times\epsilon^{\mu\nu\lambda\tau}p_{\lambda}\,\Delta_{\tau\kappa}(\bar{y})\left[\partial_{y}^{\kappa}\frac{n_{\nu}(y)}{|p\cdot n(y)|}\Delta\zeta(y,x)\right]\Bigg|_{q=0,y=\bar{y}(x,p)}\,.$
Using eq. (S35), one can prove that:
$\epsilon^{\mu\nu\lambda\tau}p_{\lambda}\Delta_{\tau\kappa}\left.\left[\partial_{y}^{\kappa}\frac{n_{\nu}(y)}{|p\cdot n(y)|}\right]\right|_{y=\bar{y}}$ $=$ $(p^{\mu}p_{\lambda}-p^{2}g^{\mu}_{\lambda})\epsilon^{\lambda\nu\kappa\tau}\frac{n_{\tau}(\bar{y})}{p\cdot n(\bar{y})}\left.\left[\partial_{\kappa}^{y}\frac{n_{\nu}(y)}{|p\cdot n(y)|}\right]\right|_{y=\bar{y}}$ (S54)
$-\epsilon^{\lambda\kappa\tau\mu}p_{\lambda}\frac{n_{\tau}(\bar{y})}{p\cdot n(\bar{y})}p^{\nu}\left.\left[\partial_{\kappa}^{y}\frac{n_{\nu}(y)}{|p\cdot n(y)|}\right]\right|_{y=\bar{y}}\,.$
As shown in the previous section, the terms in the above equation vanish, indicating that the curvature of $\Sigma_{\rm D}$ does not contribute to ${\cal A}_{\text{LE},\zeta}^{\mu}$ in eq. (S53). On the other hand, the contribution from $\partial_{y}\zeta$ reads
${\cal A}_{\text{LE},\zeta}^{\mu}(x,p)$ $=$ $\frac{2}{(2\pi)^{3}}\sum_{\bar{y}(x,p)}n_{F}(x,p)[1-n_{F}(x,p)]\delta(p^{2}-m^{2})$ (S55)
$\times\epsilon^{\mu\nu\rho\lambda}p_{\nu}\,\frac{n_{\rho}(\bar{y})}{|p\cdot n(\bar{y})|}\partial_{\lambda}^{\bar{y}}\zeta(\bar{y})\,.$
Plugging eqs. (S22) and (S55) into eq. (22), the spin Hall effect in (22) is obtained.
S5 Isothermal decoupling hypersurface
@@
$\xi^{\mu\nu}$ $=$ $\frac{1}{2T}(\partial^{\mu}u^{\nu}+\partial^{\nu}u^{\mu})-\frac{1}{2T^{2}}(u^{\nu}\partial^{\mu}T+u^{\mu}\partial^{\nu}T)\,$, (S56)
and the gradient of $\zeta$ is split as:
$\partial^{\nu}\zeta=\frac{1}{T}\partial^{\nu}\mu-\frac{\mu}{T^{2}}\partial^{\nu}T\,.$ (S57)
Substituting them into eq. (23), the vorticity-induced polarization, shear-induced polarization, and the spin Hall effect read:
@@
$S_{\text{shear}}^{\mu}(p)$ $=$ $-\frac{1}{8mN_{p}}\int_{\Sigma_{\rm D}}\!\!{\rm d}\Sigma(x)\cdot p\,n_{F}(x,p)(1-n_{F}(x,p))\epsilon^{\mu\nu\rho\lambda}p_{\nu}$
@@
$S^{\mu}_{\rm SHE}(p)$ $=$ $\frac{1}{4mN_{p}}\int_{\Sigma_{\rm D}}\!\!{\rm d}\Sigma(x)\cdot p\,n_{F}(x,p)(1-n_{F}(x,p))\epsilon^{\mu\nu\rho\lambda}p_{\nu}\sum_{\bar{y}(x,p)}\text{sgn}[p\cdot n]\frac{n_{\rho}}{Tp\cdot n}\left[\partial_{\lambda}\mu-\frac{\mu}{T}\partial_{\lambda}T\right]\Bigg|_{\bar{y}(x,p)}\,$, (S58)
@@
$S_{\text{T-grad}}^{\mu}(p)$ $=$ $-\frac{1}{8mN_{p}}\int_{\Sigma_{\rm D}}\!\!{\rm d}\Sigma(x)\cdot p\,n_{F}(x,p)[1-n_{F}(x,p)]\epsilon^{\mu\nu\rho\lambda}p_{\nu}$ (S59)
$\times\sum_{\bar{y}(x,p)}\text{sgn}[p\cdot n]\frac{1}{T^{2}}\left\{u_{\lambda}\left[\partial_{\rho}T-n_{\rho}\frac{p\cdot\partial T}{p\cdot n}\right]-\frac{p\cdot u-2\mu}{p\cdot n}n_{\rho}\partial_{\lambda}T\right\}\Bigg|_{\bar{y}(x,p)}\,.$